OpenStudy (asanchez):

I have a rhombus, ABDC AB= 2X-3 BC=3X-10 CD=X+4 AC=18-4 Whats the value of X?

OpenStudy (anonymous):

you're missing side BD, but based on teh info you have, AB should be equal to CD: 2x-3=x+4, which makes x=7, so each side is 11, but that doesn't makes sense because AC is 18-4=14, and all sides should be equal. so, i think there is something wrong. are you sure you copied the problem right?

OpenStudy (asanchez):

wait i think i typed this wrong i will retype the question...

OpenStudy (anonymous):

ok

OpenStudy (asanchez):

Rhombus ABDC AB= 2X-3 BD= 3X-10 CD=X+4 AC=18-X Its says to find the values of the variables, then find the lengths of the sides.

OpenStudy (anonymous):

okay. so AB = DC, which means 2x-3 = x+4, so x=7 We can also see that x=7 by saying BD = AC, which means 3x-10=18-x, so x=7. In any case, x=7, and each side is 11 units long.

OpenStudy (asanchez):

THANK YOU SOOO MUCHH! (: